# Asymptotic behaviors of subcritical branching killed spectrally negative Lévy process

### v1

@@ Abstract.
In this paper, we investigate the asymptotic behaviors of the survival probability and maximal displacement of a subcritical branching killed spectrally negative Lévy process $X$ in $\mathbb{R}$. Let $\zeta$ denote the extinction time, $M_{t}$ be the maximal position of all the particles alive at time $t$, and $M:=\sup_{t\geq 0}M_{t}$ be the all-time maximum. Under the assumption that the offspring distribution satisfies the $L\log L$ condition and some conditions on the spatial motion, we find the decay rate of the survival probability $\mathbb{P}_{x}(\zeta>t)$ and the tail behavior of $M_{t}$ as $t\to\infty$. As a consequence, we establish a Yaglom-type theorem. We also find the asymptotic behavior of $\mathbb{P}_{x}(M>y)$ as $y\to\infty$.
@@
The research of Yaping Zhu is supported by the China Postdoctoral Science Foundation (No. 2024M760056).
*Yaping Zhu is the cooresponding author
@@ 1. Introduction
1.1. Background and motivation
A branching Lévy process on $\mathbb{R}$ is defined as follows: at time 00, there is a particle at $x\in\mathbb{R}$ and it moves according to a Lévy process $(\xi_{t},\mathbf{P}_{x})$ on $\mathbb{R}$. After an exponential time with parameter $\beta>0$, independent of the spatial motion, this particle dies and is replaced by $k$ offspring with probability $p_{k}$, $k\geq 0$. The offspring move independently according to the same Lévy process from their birthplace and obey the same branching mechanism as their parent. This procedure goes on. Let $N_{t}$ be the set of particles alive at time $t$ and for each $u\in N_{t}$, we denote by $X_{u}(t)$ the position of $u$ at time $t$. Also, for any $u\in N_{t}$ and $s\leq t$, we use $X_{u}(s)$ to denote the position of $u$ or its ancestor at time $s$. Then the point process $Z=(Z_{t})_{t\geq 0}$ defined by
@@ 1.1. Background and motivation
be the extinction time of $Z$. Note that $\widetilde{\zeta}$ is equal in law to that of the extinction time of a continuous-time Galton-Waston process with the same branching mechanism as the branching Lévy process. Let $m:=\sum_{k=0}^{\infty}kp_{k}$ be the mean number of offspring. It is well-known that $Z$ will become extinct in finite time with probability 1 if and only if $m<1$ (subcritical) or $m=1$ and $p_{1}\neq 1$ (critical). Moreover, the process $Z$ survives with positive probability when $m>1$ (supercritical).
In the critical case, i.e. $m=1$ and $p_{1}\neq 1$, Sawyer and Fleischman [15] investigated the tail behavior of the all time maximal position of branching Brownian motion under the assumption that the offspring distribution has finite third moment. For a critical branching random walk with spatial motion having finite $(4+\varepsilon)$th moment, the tail behavior of the all time maximum was obtained by Lalley and Shao [12]. Hou et al. [8] studied the asymptotic behavior of the all time maximum of critical branching Lévy processes with offspring distribution belonging to the domain of attraction of an $\alpha$-stable distribution with $\alpha\in(1,2]$, under some assumptions on the spatial motion.
In the subcritical case, Profeta [14] gave the asymptotic behavior of the all time maximal position under the assumption that the offspring distribution has finite third moment. For related results about subcritical branching random walks, we refer the reader to [13].
@@
In the critical case, i.e., when $m=1$ and $p_{1}\neq 1$, the asymptotic behaviors of the tails of the extinction time and the maximal displacement of $Z^{0}$ were established in [9] under the assumption that the offspring distribution belongs to the domain of attraction of an $\alpha$-stable distribution, $\alpha\in(1,2]$, and some moment assumptions on the spatial motion. They also proved that the scaling limit under $\mathbb{P}_{\sqrt{t}y}(\cdot|\zeta>t)$ can be represented in terms of a super killed Brownian motion. In the subcritical case, i.e., $m\in(0,1)$, under the assumption $\sum^{\infty}_{k=1}k(\log k)p_{k}<\infty$, the asymptotic behaviors of the survival probability and the all time maximal position of branching killed Brownian motion with drift were established in [10] recently.
The purpose of this paper is to extend the results of [10] to subcritical branching killed spectrally negative Lévy processes. This extension is quite challenging since properties of Brownian motion were used crucially in [10]. Fluctuation theory of spectrally negative Lévy processes will play an important role in this paper. Another important tool is the conditioned limit theorem in Theorem 3.5 below.
@@ 1.2. Main results
Before we state our main results, we introduce some notation and some basic results on spectrally negative Lévy processes. We always assume that the offspring distribution is subcritical, i.e., $m\in(0,1)$. Let $\alpha:=\beta(1-m)$ and let $f$ be the generating function of the offspring distribution, i.e. $f(s)=\sum_{k=0}^{\infty}p_{k}s^{k}$, $s\in[0,1]$. Define
@@
where $\varphi(u)=\frac{\Phi(u)-\alpha u}{u}$ for $u\in(0,1]$ and $\varphi(0)=\Phi^{\prime}(0+)-\alpha=0$. According to [10, Lemma 2.7], $\varphi(\cdot)$ is increasing on $[0,1]$ and under the condition
@@
In this paper, we always assume that $\xi=((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a Lévy process on $\mathbb{R}$ and we will use $\mathbf{E}_{x}$ to denote expectation with respect to $\mathbf{P}_{x}$. For any $z\in\mathbb{R}$, define
@@
It follows from [9, Lemma 2.8] that, if $\mathbf{E}_{0}(\xi_{1})=0$ and $\mathbf{E}_{0}(\xi_{1}^{2})\in(0,\infty)$, then $\mathbf{E}_{x}|\xi_{\tau_{0}^{-}}|<\infty$ and the function
(1.16) $R(x):=x-\mathbf{E}_{x}\left(\xi_{\tau_{0}^{-}}\right)=-\mathbf{E}_{0}\left(\xi_{\tau_{-x}^{-}}\right),\quad x\geq 0,$
satisfies the following:
@@
Now we introduce some basics on spectrally negative Lévy processes. Let $\xi$ be a spectrally negative Lévy process on $\mathbb{R}$. For any $\lambda\in\mathbb{C}$ with $\mathrm{Re}(\lambda)\geq 0$, the Laplace exponent of $\xi$ is given by
(1.19) $\Psi(\lambda)=\ln\mathbf{E}_{0}\left[e^{\lambda\xi_{1}}\right]=a\lambda+\frac{\eta^{2}}{2}\lambda^{2}+\int_{-\infty}^{0}\left(e^{\lambda x}-1-\lambda x1_{\{|x|<1\}}\right)\Pi(\mathrm{d}x),$
where $a\in\mathbb{R}$, $\eta\geq 0$ and the Lévy measure $\Pi$ satisfies $\int_{-\infty}^{0}\left(1\land x^{2}\right)\Pi(\mathrm{d}x)<+\infty$. It is well known that
$\mathbf{E}_{0}\left[\xi_{1}\right]=\Psi^{\prime}(0+)=a+\int_{(-\infty,-1]}x\Pi(\mathrm{d}x)\in[-\infty,\infty).$
@@
(H2) The law of $\xi_{1}$ is non-lattice.
For $q\geq 0$, let
(1.20) $\psi(q):=\sup\{\lambda\geq 0:\Psi(\lambda)=q\}$
be the right inverse of $\Psi$. By Kyprianou [11, Theorem 8.1], for any $q\geq 0$, there exists a scale function $W^{(q)}:\mathbb{R}\to[0,\infty)$ such that $W^{(q)}(x)=0$ for $x<0$ and $W^{(q)}$ is a strictly increasing and continuous function on $[0,\infty)$ with Laplace transform
In the case when $\xi$ is a standard Brownian motion with drift $-b$, by using tables of Laplace transforms, one can easily get that
For any $c\geq 0$ and $x\in\mathbb{R}$, since $\{e^{c(\xi_{t}-x)-\Psi(c)t}:t\geq 0\}$ is a $\mathbf{P}_{x}$-martingale, we can define the change of measure
@@
where $\mathcal{F}_{t}:={\sigma\{\xi_{s}:s\leq t\}}$, $t\geq 0$. According to [11, Theorem 3.9], $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ is also a spectrally negative Lévy process and its Laplace exponent $\Psi_{c}(\lambda)$ is given by $\Psi_{c}(\lambda):=\Psi(\lambda+c)-\Psi(c)$. We will use $\mathbf{E}^{c}_{x}$ to denote expectation with respect to $\mathbf{P}_{x}^{c}$. Let $W_{c}^{(q)}$ be the scale function of $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$.
According to [11, Theorem 3.9], the Lévy measure of $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ is given by $1_{(-\infty,0)}e^{cx}\Pi(\mathrm{d}x)$. Thus, for any $p\geq 1$, if $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ has finite $p$-th moment, then $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ also has finite $p$-th moment. In case $c>0$, $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ will have finite exponential moment. It is also easy to see that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ is non-lattice if and only if $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is non-lattice. In particular, if $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ satisfies (H1) (or (H2)), then $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ also satisfies (H1) (or (H2)). It follows from [11, Theorem 3.9] that, if $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a Brownian motion with drift, $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ is also a Brownian motion with drift.
When $\Psi^{\prime}(0+)<0$, we have $\psi(0)>0$. In this case, $\Psi$ is strictly convex and admits a unique minimum at a $\lambda_{*}>0$ and $\Psi(\lambda_{*})<0$, $\Psi^{\prime}(\lambda_{*})=0$ and $\Psi^{\prime\prime}(\lambda_{*})>0$. Consequently, the function $\psi$ is well-defined on $[-q_{*},+\infty)$, where $-q_{*}=\Psi(\lambda_{*})$. Taking $c=\lambda_{*}$, we define the change of measure,
@@
Then $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{\lambda_{*}})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process and its Laplace exponent is given by $\Psi_{\lambda_{*}}(\lambda):=\Psi(\lambda+\lambda_{*})-\Psi(\lambda_{*})$. It is easy to see that $\Psi_{\lambda_{*}}^{\prime}(0+)=\Psi^{\prime}(\lambda_{*})=0$. Let $\mathbf{E}_{x}^{\lambda_{*}}$ be the expectation with respect to $\mathbf{P}_{x}^{\lambda_{*}}$.
Suppose that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ has finite second moment, then $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{\lambda_{*}})_{x\in\mathbb{R}})$ also has finite second moment. Since $\mathbf{E}_{0}^{\lambda_{*}}(\xi_{1})=\Psi_{\lambda_{*}}^{\prime}(0+)=0$, by [9, Lemma 2.8], we have $\mathbf{E}_{x}^{\lambda_{*}}|\xi_{\tau_{0}^{-}}|<\infty$. Define
@@
Denote $\mathbb{R}_{+}=[0,\infty)$. Our first main result is on the large-time asymptotics of the survival probability.
Theorem 1.1.
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]=0$, then
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]>0$, then
    (1.29) $\lim_{t\to\infty}e^{\alpha t}\mathbb{P}_{x}(\zeta>t)=C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x).$
@@
    Assume further that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ satisfies (H2). If $\mathbf{E}_{0}\left[\xi_{1}\right]<0$, then
@@
Combining the result above with (1.18) and (1.22), we immediately recover the first two conclusions of [10, Theorem 1.1] as a corollary. Furthermore, when $\xi$ is a standard Brownian motion with drift $-\rho$, we have $\Psi(\lambda)=-\rho\lambda+\frac{1}{2}\lambda^{2}$ and $\lambda_{*}=\rho$. When $\rho>0$, a straightforward calculation yields that
@@
This result is consistent with [10, Theorem 1.1, (iii)], where $C_{0}(\rho)=C_{0}$.
@@
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]=0$, then for any $y\geq 0$, we have
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]>0$, then for any $y\in\mathbb{R}$, we have
    (1.33) $\lim_{t\to\infty}e^{\alpha t}\mathbb{P}_{x}\left(M_{t}>\sqrt{t}y+\mathbf{E}_{0}\left[\xi_{1}\right]t\right)=\frac{C_{sub}\mathbf{E}_{0}\left[\xi_{1}\right]W^{(0)}(x)}{\sqrt{2\pi}}\int_{\frac{y}{\sigma}}^{\infty}e^{-\frac{z^{2}}{2}}\mathrm{d}z.$
@@
    Assume further that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ satisfies (H2). If $\mathbf{E}_{0}\left[\xi_{1}\right]<0$, then for any $y\geq 0$, we have
@@
Noticing that $\mathbb{P}_{z}(M_{N}>0)=\mathbb{P}_{z}(\zeta>N)$ for $z>0$, $C_{0}$ in Theorem 1.1 and $C_{1}(0)$ in Theorem 1.2 are the same. Combining the result above with (1.18) and (1.22), we immediately recover [10, Theorem 1.3] as a corollary.
@@
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]=0$, then we have
@@
    If $\mathbf{E}_{0}\left[\xi_{1}\right]>0$, then we have
    (1.36) $\mathbb{P}_{x}\left(\frac{M_{t}-\mathbf{E}_{0}\left[\xi_{1}\right]t}{\sqrt{t}}\in\cdot\Big{|}\zeta>t\right)\quad\stackrel{{\scriptstyle\mathrm{d}}}{{\Longrightarrow}}\quad N(0,\sigma^{2}),$
    where $N(0,\sigma^{2})$ is normal distribution with mean 00 and variance $\sigma^{2}$.
@@
    Assume further that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ satisfies (H2). If $\mathbf{E}_{0}\left[\xi_{1}\right]<0$, then there exists a random variable $(X,\mathbb{P})$ whose law is independent of $x$ such that
@@
Our third main result is on the asymptotic behavior of the all-time maximum $M$. Note that we do not need any assumption on the spectrally Lévy process in the following result.
@@
Assume that (1.7) holds and that $((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process. There exits a constant $C_{2}(\alpha)\in(0,1]$ such that for any $x>0$,

### v2

# Asymptotic behaviors of subcritical branching killed Lévy processes

@@ Abstract.
In this paper, we investigate the asymptotic behaviors of the survival probability and maximal displacement of a subcritical branching killed Lévy process $X$ in $\mathbb{R}$. Let $\zeta$ denote the extinction time, $M_{t}$ be the maximal position of all the particles alive at time $t$, and $M:=\sup_{t\geq 0}M_{t}$ be the all-time maximum. Under the assumption that the offspring distribution satisfies the $L\log L$ condition and some conditions on the spatial motion, we find the decay rate of the survival probability $\mathbb{P}_{x}(\zeta>t)$ and the tail behavior of $M_{t}$ as $t\to\infty$. As a consequence, we establish a Yaglom-type theorem. We also find the asymptotic behavior of $\mathbb{P}_{x}(M>y)$ as $y\to\infty$.
@@
The research of Yaping Zhu is supported by the Fundamental Research Funds for the Central Universities.
*Yaping Zhu is the corresponding author
@@ 1. Introduction
$1.$ Background and motivation
A branching Lévy process on $\mathbb{R}$ is defined as follows: at time 0, there is a particle at $x\in\mathbb{R}$ and it moves according to a Lévy process $(\xi_{t},\mathbf{P}_{x})$ on $\mathbb{R}$. After an exponential time with parameter $\beta>0$, independent of the spatial motion, this particle dies and is replaced by $k$ offspring with probability $p_{k}$, $k\geq 0$. The offspring move independently according to the same Lévy process starting from the death position of their parent. This procedure goes on. Let $N_{t}$ be the set of particles alive at time $t$ and for each $u\in N_{t}$, we denote by $X_{u}(t)$ the position of $u$ at time $t$. Also, for any $u\in N_{t}$ and $s\leq t$, we use $X_{u}(s)$ to denote the position of $u$ or its ancestor at time $s$. Then the point process $Z=(Z_{t})_{t\geq 0}$ defined by
@@
be the extinction time of $Z$. Note that $\widetilde{\zeta}$ is equal in law to that of the extinction time of a continuous-time Galton-Watson process with the same branching mechanism as the branching Lévy process. Let $m:=\sum_{k=0}^{\infty}kp_{k}$ be the mean number of offspring. It is well-known that $Z$ will become extinct in finite time with probability 1 if and only if $m<1$ (subcritical) or $m=1$ and $p_{1}\neq 1$ (critical). Moreover, the process $Z$ survives with positive probability when $m>1$ (supercritical).
@@
In the critical case, i.e., when $m=1$ and $p_{1}\neq 1$, the asymptotic behaviors of the tails of the extinction time and the maximal displacement of $Z^{0}$ were established in [11] under the assumption that the offspring distribution belongs to the domain of attraction of an $\alpha$-stable distribution, $\alpha\in(1,2]$, and some moment assumptions on the spatial motion. It was also shown in [11] that the scaling limit under $\mathbb{P}_{\sqrt{t}y}(\cdot|\zeta>t)$ can be represented in terms of a super killed Brownian motion. In the subcritical case, i.e., $m\in(0,1)$, under the assumption $\sum^{\infty}_{k=1}k(\log k)p_{k}<\infty$, the asymptotic behaviors of the survival probability and the all time maximal position of branching killed Brownian motion with drift were established in [12] recently.
The asymptotic behavior of branching Lévy processes have been studied earlier. In the critical case, i.e. $m=1$ and $p_{1}\neq 1$, Sawyer and Fleischman [18] investigated the tail behavior of the all time maximal position of branching Brownian motion under the assumption that the offspring distribution has finite third moment. For a critical branching random walk with spatial motion having finite $(4+\varepsilon)$th moment, the tail behavior of the all time maximum was obtained by Lalley and Shao [14]. Hou et al. [10] studied the asymptotic behavior of the all time maximum of critical branching Lévy processes with offspring distribution belonging to the domain of attraction of an $\alpha$-stable distribution with $\alpha\in(1,2]$, under some assumptions on the spatial motion. In the subcritical case, Profeta [17] gave the asymptotic behavior of the all time maximal position under the assumption that the offspring distribution has finite third moment. For related results about subcritical branching random walks, we refer the reader to [16].
The purpose of this paper is to extend the results of [12] to subcritical branching killed Lévy processes. This extension is quite challenging since properties of Brownian motion were used crucially in [12]. Fluctuation theory of Lévy processes will play an important role in this paper. Another important tool is the conditioned limit theorem in Theorem 3.5 below.
@@ 1.2. Main results
Before we state our main results, we introduce some notation and some basic results on Lévy processes. We always assume that the offspring distribution is subcritical, i.e., $m\in(0,1)$. Let $\alpha:=\beta(1-m)$ and let $f$ be the generating function of the offspring distribution, i.e. $f(s)=\sum_{k=0}^{\infty}p_{k}s^{k}$, $s\in[0,1]$. Define
@@
where $\varphi(u)=\frac{\Phi(u)-\alpha u}{u}$ for $u\in(0,1]$ and $\varphi(0)=\Phi^{\prime}(0+)-\alpha=0$. According to [12, Lemma 2.7], $\varphi(\cdot)$ is increasing on $[0,1]$ and under the condition
@@
In this paper, we always assume that $\xi=((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a Lévy process on $\mathbb{R}$ with
$\log\mathbf{E}_{x}\left(e^{i\theta(\xi_{1}-\xi_{0})}\right)=ia\theta+\frac{1}{2}\eta^{2}\theta^{2}+\int^{\infty}_{-\infty}(1-e^{i\theta x}+i\theta x1_{\{|x|<1\}})\Pi(dx),\quad\theta\in\mathbb{R},$
where $\mathbf{E}_{x}$ stands for the expectation with respect to $\mathbf{P}_{x}$, $a\in\mathbb{R}$, $\eta\geq 0$ and the Lévy measure $\Pi$ satisfies $\int_{\mathbb{R}}\left(1\land x^{2}\right)\Pi(\mathrm{d}x)<+\infty$. For any $z\in\mathbb{R}$, define
@@
Define the function
(1.16) $R(x):=x-\mathbf{E}_{x}\left(\xi_{\tau_{0}^{-}}\right)=-\mathbf{E}_{0}\left(\xi_{\tau_{-x}^{-}}\right),\quad x\geq 0.$
It follows from [11, Lemma 2.8] that if $\mathbf{E}_{0}(\xi_{1})=0$ and $\mathbf{E}_{0}(\xi_{1}^{2})\in(0,\infty)$, then $\mathbf{E}_{x}|\xi_{\tau_{0}^{-}}|<\infty$ and $R(x)$ satisfies the following:
@@
(H2) The law of $\xi_{1}$ is non-lattice, i.e., $\mathbf{P}_{x}\left(\xi_{1}\in h\mathbb{Z}+a\right)\neq 1,\forall h>0,a\in[0,h)$.
Remark 1.
Condition (H2) will be assumed in the case $\mathbf{E}_{0}(\xi_{1})<0$. In this case, we rely on the conditioned limit theorem for random walks established by [8], which requires the non-lattice condition.
In the case $\mathbf{E}_{0}\left(\xi_{1}\right)<0$, we will perform an Esscher transform on the Lévy process. For this, we assume that
(H3) The Laplace exponent $\Psi(\lambda):=\log\mathbf{E}_{0}\left(e^{\lambda\xi_{1}}\right)$ is finite for all $\lambda\in(\Lambda_{1},\Lambda_{2})$ with $\Lambda_{1}\in[-\infty,0]$ and $\Lambda_{2}\in(0,\infty]$. Moreover, there exists a unique $\lambda_{*}\in(0,\Lambda_{2})$ such that $\Psi^{\prime}(\lambda_{*})=0$.
Note that $\Psi(\lambda)$ is finite if and only if $\int_{\{|x|\geq 1\}}e^{\lambda x}\Pi(dx)<\infty$ and that for any $\lambda\in(\Lambda_{1},\Lambda_{2})$,
(1.19) $\Psi(\lambda)=a\lambda+\frac{\eta^{2}}{2}\lambda^{2}+\int_{\mathbb{R}}\left(e^{\lambda x}-1-\lambda x1_{\{|x|<1\}}\right)\Pi(\mathrm{d}x).$
Note also that $\Psi$ is convex in $(\Lambda_{1},\Lambda_{2})$.
Remark 2.
If $\xi=((\xi_{t})_{t\geq 0},(\mathbf{P}_{x})_{x\in\mathbb{R}})$ is a spectrally negative Lévy process, then $\xi$ has finite Laplace exponent in $(0,\infty)$. If $\xi$ is a spectrally negative Lévy process with $\mathbf{E}_{0}\left(\xi_{1}\right)<0$, then $\Psi$ admits a unique minimum at a $\lambda_{*}>0$ and $\Psi(\lambda_{*})<0$, $\Psi^{\prime}(\lambda_{*})=0$ and $\Psi^{\prime\prime}(\lambda_{*})>0$. So in this case (H3) is automatically satisfied.
For any $c\in(\Lambda_{1},\Lambda_{2})$ and $x\in\mathbb{R}$, since $\{e^{c(\xi_{t}-x)-\Psi(c)t}:t\geq 0\}$ is a $\mathbf{P}_{x}$-martingale, we can define the change of measure
@@
where $\mathcal{F}_{t}:={\sigma\{\xi_{s}:s\leq t\}}$, $t\geq 0$. According to [13, Theorem 3.9], $\xi^{(c)}=((\xi_{t})_{t\geq 0},(\mathbf{P}_{x}^{c})_{x\in\mathbb{R}})$ is also a Lévy process and its Laplace exponent $\Psi_{c}(\lambda)$ is given by $\Psi_{c}(\lambda):=\Psi(\lambda+c)-\Psi(c)$. We will use $\mathbf{E}^{c}_{x}$ to denote expectation with respect to $\mathbf{P}_{x}^{c}$.
Recall that $\int_{\{|x|\geq 1\}}e^{\lambda x}\Pi(dx)<\infty$ for any $\lambda\in(\Lambda_{1},\Lambda_{2})$. According to [13, Theorem 3.9], the Lévy measure of $\xi^{(c)}$ is given by $e^{cx}\Pi(\mathrm{d}x)$. Combining the two facts above, we get that, if $\xi$ has finite $p$-th moment with $p\geq 1$, then for any $c\in(\Lambda_{1},\Lambda_{2})$, $\xi^{(c)}$ also has finite $p$-th moment and so $\xi^{(c)}$ satisfies (H1). It is also easy to see that $\xi^{(c)}$ is non-lattice if and only if $\xi$ is non-lattice. We note that, by [13, Theorem 3.9], (i) if $\xi$ is a spectrally negative Lévy process with Laplace exponent $\Psi$, then $\xi^{(c)}$ is a spectrally negative Lévy process with Laplace exponent $\Psi_{c}(\lambda)$ given by $\Psi_{c}(\lambda):=\Psi(\lambda+c)-\Psi(c)$; and (ii) if $\xi$ is a Brownian motion with drift, $\xi^{(c)}$ is also a Brownian motion with drift.
When $\mathbf{E}_{0}\left(\xi_{1}\right)<0$ and (H3) holds, we take $c=\lambda_{*}$ and define the change of measure
@@
Then $\xi^{(\lambda_{*})}$ is a Lévy process and its Laplace exponent is given by $\Psi_{\lambda_{*}}(\lambda):=\Psi(\lambda+\lambda_{*})-\Psi(\lambda_{*})$. It is easy to see that $\Psi_{\lambda_{*}}^{\prime}(0+)=\Psi^{\prime}(\lambda_{*})=0$. Let $\mathbf{E}_{x}^{\lambda_{*}}$ be the expectation with respect to $\mathbf{P}_{x}^{\lambda_{*}}$. If $\xi$ satisfies (H1), then since $\mathbf{E}_{0}^{\lambda_{*}}(\xi_{1})=\Psi_{\lambda_{*}}^{\prime}(0+)=0$, by [11, Lemma 2.8], we have $\mathbf{E}_{x}^{\lambda_{*}}|\xi_{\tau_{0}^{-}}|<\infty$. Define
@@
Denote $\mathbb{R}_{+}=[0,\infty)$. Let $\sigma^{2}:=\mathbf{E}_{0}(\xi_{1}^{2})$. Our first main result is on the large-time asymptotic behavior of the survival probability.
Theorem $1.$
Assume (1.7) holds and $\xi$ is a Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)=0$, then
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)>0$, then
    (1.26) $\lim_{t\to\infty}e^{\alpha t}\mathbb{P}_{x}(\zeta>t)=q_{x}C_{sub},$
    where $q_{x}:=\mathbf{P}_{x}\left(\tau_{0}^{-}=\infty\right)>0.$
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)<0$ and $\xi$ satisfies (H2) and (H3), then
@@
[12, Theorem 1.1] investigates the asymptotic behavior of the survival probability of a branching killed Brownian motion with drift $\rho$. The first two statements of [12, Theorem 1.1] are as follows.
(1) if $\rho=0$, then $\lim_{t\to\infty}\sqrt{t}e^{\alpha t}\mathbf{P}_{x}(\zeta>t)=\sqrt{\frac{2}{\pi}}C_{sub}x.$
(2) If $\rho<0$, then $\lim_{t\to\infty}e^{\alpha t}\mathbf{P}_{x}(\zeta>t)=\left(1-e^{2\rho x}\right).$
Combining Theorem 1.1 (1) and (2) with (1.18), we immediately recover the first two conclusions of [12, Theorem 1.1]. Furthermore, when $\xi$ is a standard Brownian motion with drift $\rho$, we have $\Psi(\lambda)=-\rho\lambda+\frac{1}{2}\lambda^{2}$ and $\lambda_{*}=\rho$. When $\rho>0$, a straightforward calculation yields that
@@
This result is consistent with [12, Theorem 1.1, (iii)], where $C_{0}(\rho)=C_{0}$.
@@
Assume (1.7) holds and $\xi$ is a Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)=0$, then for any $y\geq 0$, we have
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)>0$, then for any $y\in\mathbb{R}$, we have
    (1.30) $\lim_{t\to\infty}e^{\alpha t}\mathbb{P}_{x}\left(M_{t}>\sqrt{t}y+\mathbf{E}_{0}\left(\xi_{1}\right)t\right)=\frac{q_{x}C_{sub}}{\sqrt{2\pi}}\int_{\frac{y}{\sigma}}^{\infty}e^{-\frac{z^{2}}{2}}\mathrm{d}z.$
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)<0$ and $\xi$ satisfies (H2) and (H3), then for any $y\geq 0$, we have
@@
Note that $\mathbb{P}_{z}(M_{N}>0)=\mathbb{P}_{z}(\zeta>N)$ for $z>0$, thus $C_{0}$ in Theorem 1.1 and $C_{1}(0)$ in Theorem 1.2 are the same. Combining the result above with (1.18) and (1.37), we immediately recover [12, Theorem 1.3] as a corollary.
@@
Assume (1.7) holds and $\xi$ is a Lévy process satisfying (H1). Let $x>0$.
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)=0$, then we have
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)>0$, then we have
    (1.33) $\mathbb{P}_{x}\left(\frac{M_{t}-\mathbf{E}_{0}\left(\xi_{1}\right)t}{\sqrt{t}}\in\cdot\Big|\zeta>t\right)\quad\stackrel{{\scriptstyle\mathrm{d}}}{{\Longrightarrow}}\quad N(0,\sigma^{2}),$
    where $N(0,\sigma^{2})$ is normal distribution with mean 0 and variance $\sigma^{2}$.
@@
    If $\mathbf{E}_{0}\left(\xi_{1}\right)<0$ and $\xi$ satisfies (H2) and (H3), then there exists a random variable $(X,\mathbb{P})$ whose law is independent of $x$ such that
@@
In the following theorem we assume that $\xi$ is a spectrally negative Lévy process with Laplace exponent $\Psi$. For $q\geq 0$, let
(1.35) $\psi(q):=\sup\{\lambda\geq 0:\Psi(\lambda)=q\}$
be the right inverse of $\Psi$. By Kyprianou [13, Theorem 8.1], for any $q\geq 0$, there exists a scale function $W^{(q)}:\mathbb{R}\to[0,\infty)$ such that $W^{(q)}(x)=0$ for $x<0$ and $W^{(q)}$ is a strictly increasing and continuous function on $0,\infty)$ with Laplace transform
In the case when $\xi$ is a standard Brownian motion with drift b-b, by using tables of Laplace transforms, one can easily get that
Our third main result is on the asymptotic behavior of the all-time maximum $M$ of branching killed spectrally negative Lévy process.
@@
Assume that (1.7) holds and that $\xi$ is a spectrally negative Lévy process. There exists a constant $C_{2}(\alpha)\in(0,1]$ such that for any $x>0$,